A couple just had a baby. How much should they invest now at 5.5% compounded weekly in order to have $60,000 for the child’s education 19 years from now? Round to the nearest dollar.
Answer by nerdybill(7384)

A couple just had a baby. How much should they invest now at 5.5% compounded weekly in order to have $60,000 for the child’s education 19 years from now? Round to the nearest dollar.
.
Compound interest formula:
A = P(1+r/n)^(nt)
where
A=60000
P is what we're looking for
r=.055
n=52
t=19
.
plug it all in
60000 = P(1+.055/52)^(52*19)
60000 = P(1+0.00105769231)^988
60000 = P(1.00105769231)^988
60000 = P(2.8418286765312691531872438722308)
60000/2.8418286765312691531872438722308 = P
$21113.17 = P